# RiskLabs: Predicting Financial Risk Using Large Language Model Based on Multi-Sources Data

### v1

@@ 2.5 News-Market-Reactions Encoder
Figure 3: This figure illustrates the pipeline for enriching the news information. First, the pipeline will analyze the sentiments from the target news. Then, based on the binary questions bank designed for different topics, Pipeline can extract the information and answer these questions. Finally, if the pipeline could capture the signal of a specific topic, it would also give feedback on the potential market response.
Initially, we collect news pertaining to a particular company from the three days prior to our target trading day, including the market responses for each day. This compilation allows a Large Language Model (LLM) to examine the data and deduce how the news might influence the stock’s performance in the subsequent days. Subsequently, our strategy involves identifying historical dates on which the news profiles closely resemble those on the day before the intended trading day. This comparative analysis aims to understand the potential market reactions based on similar past events, thereby informing our trading strategies with a nuanced perspective on news impact
Based on these news features, we propose a news-market reactions encoder composed of two main components. (See Algorithm 1) Firstly, we collect news for a specific company from the three days prior to our target trading day, along with the market reactions for each days. We then use an LLM to analyze the data and identify how the news might affect the stock’s performance in the days that follow. Secondly, we attempt to identify the historical date whose news is similar to that of the intended trading date. In this scenario, we face the challenge of assessing similarity once more. Given that we are comparing two groups of news, the likelihood of identical combinations of news reoccurring is extremely small, which may lead to less accurate retrieval results. (See Figure 3)
@@
Algorithm 1 Enriched News Pipeline
1:procedure ExtractNewsAttributes($newsArticle$)
[…]
4:     if sentiment == positive then
5:         print "Sentiment towards the company is positive."
6:     else if sentiment == neutral then
7:         print "Sentiment towards the company is neutral."
8:     else
9:         print "Sentiment towards the company is negative."
10:     end if
[…]
12:     if financialMetrics mentioned then
[…]
14:         print "Financial trend is", financialTrend "."
15:     end if
[…]
17:     if regulatoryIssues mentioned then
[…]
19:         print "Regulatory issue category:", complianceCategory "."
20:     end if
21:end procedure
After processing all news in our database through this pipeline, we assign these identified attributes to each news group as metadata. This metadata becomes instrumental in locating news similar to those of the target trading days. Rather than directly comparing two groups of news, we initially compare the metadata between them. This process helps us identify the top k news groups in the historical record that share similar metadata attributes. Subsequently, we assess the similarity between these top k groups and the news groups from the target trading date. (See Figure 4)
Figure 4: This diagram illustrates the process by which the News Analyzer assesses similarities across various news collections. As news items pass through the enrichment pipeline, they are tagged with multiple attributes. Each circle in the figure represents one of these attributes. To identify similar news collections from historical data, the analyzer starts by comparing these attributes and filtering out certain ones. Subsequently, it evaluates the similarity among the remaining attributes to determine the connections between different news collections.
@@
In this way, it is efficient to identify the two groups that are similar, then we can make an inference for the target trading date about how these results could impact the following stock market based on the historical similar news and its corresponding market reactions.
2.6 Multi-Task Prediction
We start the prediction by aggregating features from various modules to form a comprehensive feature representation. This unified representation is then connected to a two-layer neural network, which is employed to carry out the regression task. This integration is essential for synthesizing the diverse inputs into a cohesive output, leveraging the strengths of each module for effective analysis and prediction.
Building on insights from previous research in multimodal financial risk prediction, which has demonstrated substantial improvements in prediction performance through multitask learning, we adopt a joint modeling approach. Here, we concurrently model volatility prediction and VaR prediction using a multi-task framework. The multi-task prediction module is comprised of two separate single-layer feedforward networks, each responsible for predicting volatility (vol) and Value at Risk (var) values individually. We train RiskLabs by optimizing multitask loss:
$\mathcal{L}=\mu(\sum_{i}(\hat{y_{i}}-y_{i})^{2})+(1-\mu)\text{max}(q\times(v-\hat{v}),(1-q)(\hat{v}-v)).$ (14)
2.7 Time Decay Hyper-parameter and Dynamic Moving Time Window Training
@@
Figure 5: This figure illustrates the methodology for modeling the impact of earnings conference calls on the stock market when earnings conference calls are not available. It uses a curve to represent the time decay effect of an earnings conference call’s influence over time. When an earnings conference call is initially issued, its impact on the market is at its peak. Over time, this influence gradually diminishes.
@@
Figure 6: This figure provides a visual representation of the Rolling Window methodology in action. By establishing a fixed window length, the approach systematically progresses the training set forward, day by day, in alignment with the passage of time.
3 Experiment and Results
In this section, we aim to thoroughly evaluate the effectiveness of RiskLabs by conducting extensive experiments. Our approach begins with outlining the comparative methodology we have selected, along with the specifics of our implementation details. This is followed by a detailed comparison of RiskLabs’ performance against other existing methods, providing a clear understanding of its relative efficacy. Then, to gain deeper insights into the contribution of each component within RiskLabs, we conduct a series of ablation studies. These experiments are designed to dissect and analyze the function and impact of each module individually, offering a granular view of how each part contributes to the overall performance of the framework.
3.1 Experiment Setup
3.1.1 Baselines
@@
  • Classical Methods: The analysis incorporates the GARCH model, a classical auto-regressive model for predicting volatility, as well as its various derivatives Franses and Van Dijk (1996); Kim and Won (2018). These models are widely recognized and commonly employed in the realm of volatility prediction. Primarily, they are tailored for short-term volatility forecasting and may not perform as efficiently in predicting average volatility over an extended period (e.g., n-day volatility).
  • LSTM Gers et al. (2000): Long Short-Term Memory Networks (LSTMs) are a popular choice for financial time series prediction due to their efficacy in handling sequential data. In the context of volatility prediction, we select a straightforward LSTM model as a benchmark.
  • MT-LSTM+ATT Luong et al. (2015): merges the prediction of average n-day volatility with the forecasting of single-day volatility, using attention-enhanced LSTM as the foundational learning models.
  • HAN (Glove): This baseline implements a Hierarchical Attention Network that incorporates dual-layered attention mechanisms at both the word and sentence levels. Initially, every word within a sentence is transformed into a word embedding through the pre-trained Glove 300-dimensional embeddings. Subsequently, these embedded sentences are processed by a Bi-GRU encoder Chung et al. (2014), and in parallel, another Bi-GRU encoder is utilized to formulate a representation of each document as a series of sentences. This representation of the document is then fed into the concluding regression layer to generate predictions.
  • MRDM Qin and Yang (2019): The MRDM model first introduced a multi-modal deep regression approach for volatility prediction tasks. It utilizes pre-trained GloVe embeddings and bespoke acoustic features, which are processed through individual BiLSTMs to generate uni-modal contextual embeddings. These embeddings are subsequently merged and input into a two-layer dense network for further processing.
  • HTML Yang et al. (2020): This work presented a state-of-the-art model that employs WWM-BERT for text token encoding. Similar to MDRM, HTML also leverages the same audio features. These unimodal features are then combined and processed through a sentence-level transformer, resulting in multimodal representations for each call.
@@
3.1.2 Implementation Detail
We use GPT-4 for the Earnings Conference Call Analyzer and the News-Market Reactions Encoder, utilizing it to analyze ECC data and build the news database. Throughout this process, we set the temperature parameter to 0. This ensures that the Large Language Models (LLMs) produce the most predictable responses, which aids in maintaining consistency in our experiments.
For the overall training of the framework, we developed the code using PyTorch. Each Multi-Head Attention layer in the network comprises 8 individual heads, and the training process utilized batch sizes $b\in\{2,4,8\}$. We use a grid search to determine the optimal parameters and select the learning rate $\lambda$ for Adam optimizer among $\{1e-3,1e-5,1e-6,1e-7\}$. The best hyper-parameters were kept consistent across all experiments, with the exception of the trade-off parameter $\mu$ which varied between the two tasks.
3.2 Performance Comparison (RQ1 & RQ2 & RQ3)
Table 3: Performance results on our proposed framework RiskLabs from different baseline models.
Model $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$ $VaR$ Multi-Task
Classical Method 0.713 1.710 0.526 0.330 0.284 / $\otimes$
LSTM 0.746 1.970 0.459 0.320 0.235 / $\otimes$
MT-LSTM-ATT 0.739 1.983 0.435 0.304 0.233 / $\otimes$
HAN 0.598 1.426 0.461 0.308 0.198 / $\otimes$
MRDM 0.577 1.371 0.420 0.300 0.217 / $\otimes$
HTML 0.401 0.845 0.349 0.251 0.158 / $\checkmark$
GPT-3.5-Turbo 2.198 2.152 1.793 2.514 2.332 0.371 $\checkmark$
RiskLabs 0.324 0.585 0.317 0.233 0.171 0.049 $\checkmark$
Table 4: Comparison of Value at Risk Predictions: AI Techniques vs. Traditional Financial Methods. The predefined VaR value is 0.05.
Method Prediction of VaR
Historical Method 0.016
Fully Connected Neural Network 0.044
LSTM 0.056
RiskLabs 0.049
The predefined VaR value is 0.05, meaning that the closer the model’s predictions are to 0.05, the better its performance. The result above shows that the prediction of VaR via applying the historical method is 0.016, which is significantly below the pre-defined percentile(5%). It indicates that the historical method overestimated the 95% VaR benchmark. Tracing back to 2016, we know there was a global financial crisis in 2015, and its effect lasted till the beginning of 2016. On January 20, 2016, the price of crude oil fell below \$27 a barrel; the DJIA index took a roller coaster from down 565 points to down 249 intraday. In February, the YTD (yield to Date) return came to -10.5%. These events, together with the sequelae of the 2015 stock market crisis, define 2016 as a risky year. Comparing 2016, 2017 will be much better. In January 2017, DJIA achieved a new historical height, landing above 20,000. The stock market experienced a boost with a 25% growth rate for DJIA, 19% for S&P 500, and 28% for Nasdaq. The market had strong confidence, and the VIX index in 2017 came to its historical lowest point. That explains the reason why the historical method may overestimate the 95% VaR benchmark, due to this method duplicating the extreme scenarios from 2016 to 2017, which leads to the extra estimation of financial risks.
Figure 7: This figure presents two plots that compare daily Value at Risk predictions (red curve) with the actual returns of the asset (black dot). It visualizes the percentage of actual returns exceeding the predicted VaR. For instance, with a predefined VaR of 0.05, observing approximately 5% of the actual returns surpassing the predicted VaR curve indicates a high degree of prediction accuracy. The left plot showcases the VaR forecast using the historical method, illustrating how this traditional technique estimates risk in relation to the asset’s actual performance. The right plot, on the other hand, using a fully connected neural network for VaR prediction, offering a modern computational approach to risk assessment.
@@ 3.2 Performance Comparison (RQ1 & RQ2 & RQ3)
3.3 Comparison Across Modules (RQ4)
In addition to overall performance, we also wanted to understand the different relative contributions of each module in the RiskLabs design decision to predictive performance. Here, we evaluate the impact of different data combinations on predictive performance to better understand the relative contributions of data sources and each sub-module. We designed the ablation study as follows:
Table 5: Ablation Study: performance results of different modules.
Module $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$ $VaR$
Audio + Text 0.373 0.645 0.362 0.280 0.204 0.131
Audio + Text + Analysis 0.357 0.627 0.335 0.267 0.199 0.057
Audio + Text + Analysis + VIX 0.324 0.585 0.317 0.233 0.171 0.049
Table 5 reveals that the ‘Audio + Text’ combination yields superior results compared to the baseline HTML model for 3-day forecasts. For longer periods, specifically 7, 15, and 30 days, RiskLabs demonstrates predictions closely aligned with HTML, while still outperforming other baseline models. It’s worth noting that with just ‘Audio + Text’, our model is more concise than HTML. This efficiency stems from our use of an advanced pre-trained model with extensive parameters, enabling the mapping of text and audio to vector representations. These vectors are then processed through a multi-head self-attention mechanism for feature extraction, underscoring the effectiveness of utilizing large-scale models in enhancing AI’s capability in risk prediction.
Further enhancement of the model was observed upon the integration of earnings conference call analysis text and time-series data, leading to incremental improvements in RiskLab’s performance. This improvement was particularly pronounced in medium to long-term forecasts (7, 15, and 30 days). This finding suggests that earnings conference calls, as critical investment references, exert more influence on short-term risk volatility. For long-term forecasts, the incorporation of diverse information sources appears to be instrumental in refining prediction accuracy. Cumulatively, these experimental results attest to the valuable contribution of each component within the RiskLabs framework, culminating in a robust predictive model.
3.4 Challenge and Solutions (RQ5)
Our experimental findings demonstrate that leveraging large language models to integrate diverse information sources can significantly enhance the capability of our framework in predicting financial risks. With this in mind, we’ve expanded our data sources to include variables that could influence market volatility, such as news disseminated through social media. As detailed in Section 2.4, we utilize LLM to collect and analyze daily financial news. Initial small-scale experiments indicate that this addition can further improve the model’s performance. However, scaling up has introduced challenges: 1) the variable quality of news sources, with some containing misinformation, potentially introduces noise into our news database and affects audio model predictions; 2) there is a need to test the module’s effectiveness across a broader data spectrum.
To address these challenges, we are implementing more nuanced steps to filter out low-quality news. Additionally, we are amassing newer data samples to enlarge our dataset, allowing for more comprehensive validation of our model’s efficacy. Furthermore, we will apply a "Dynamic Moving Time Window" and a "Time Decay Hyper-parameter" in RiskLabs (see in Section 2.4). These features will enable more flexible training and forecasting on a daily basis, thus offering investors timelier and more accurate risk assessments.
4 Conclusion
In this study, we explored the utilization of LLMs in predicting financial risks and introduced the RiskLabs framework. This innovative framework employs LLMs to systematically organize and analyze diverse financial data types and sources, augmenting deep learning models in financial risk prediction. Central to RiskLabs are specialized modules: the Earnings Conference Call Encoder, the Time-Series Encoder, and the News-Market Reactions Encoder, each designed to process specific financial data. These encoders collectively facilitate the merging of various data features for robust multi-task financial risk forecasting.
Our extensive evaluation of RiskLabs focused on its short-, medium-, and long-term predictive capabilities, specifically in forecasting market volatility and VaR metrics. The empirical findings from our study led to several key insights: 1) The RiskLabs framework demonstrates a high efficacy in predicting financial risks, confirming its potential as a valuable tool in this domain. 2) While LLMs in isolation may not yield effective financial risk predictions, their strategic application in processing relevant financial data significantly enhances the predictive power of deep learning models. 3) Ablation studies further affirm that each individual module of RiskLabs meaningfully contributes to the accuracy of the final risk predictions. Overall, our research underscores the transformative potential of LLMs in financial risk assessment, marking a significant step forward in the application of AI in finance.
We are currently enhancing the News-Market Reactions Encoder module within RiskLabs, aiming to refine its ability to utilize high-quality news information effectively. This enhancement is pivotal for augmenting the model’s performance in predicting financial risks. Concurrently, we are engaged in the acquisition of more recent and expansive datasets to facilitate a larger-scale validation of our model. A key aspect of our ongoing development is the implementation of dynamic training windows in RiskLabs. This feature is designed to increase the model’s adaptability, enabling it to deliver more prompt and precise risk-forecasting information to investors. These advancements are part of our continuous effort to advance RiskLabs, ensuring it remains at the forefront of financial risk prediction technology.

### v2

@@ 3.4.2. News-Market-Reactions Encoder
Figure 3. This figure illustrates the pipeline for enriching the news information. First, the pipeline will analyze the sentiments from the target news. Then, based on the binary questions bank designed for different topics, Pipeline can extract the information and answer these questions. Finally, if the pipeline could capture the signal of a specific topic, it would also give feedback on the potential market response.
Initially, we collect news pertaining to a particular company from the three days prior to our target trading day, including the market responses for each day. This compilation allows a Large Language Model (LLM) to examine the data and deduce how the news might influence the stock’s performance in the subsequent days. Subsequently, our strategy involves identifying historical dates on which the news profiles closely resemble those on the day before the intended trading day. This comparative analysis aims to understand the potential market reactions based on similar past events, thereby informing our trading strategies with a nuanced perspective on news impact.
Based on these news features, we propose a news-market reactions encoder composed of two main components. Firstly, we collect news for a specific company from the three days prior to our target trading day, along with the market reactions for each day. We then use an LLM to analyze the data and identify how the news might affect the stock’s performance in the days that follow. Secondly, we attempt to identify the historical date whose news is similar to that of the intended trading date. In this scenario, we face the challenge of assessing similarity once more. Given that we are comparing two groups of news, the likelihood of identical combinations of news reoccurring is extremely small, which may lead to less accurate retrieval results.
@@
After processing all news in our database through this pipeline, we assign these identified attributes to each news group as metadata. This metadata becomes instrumental in locating news similar to that of the target trading days. Rather than directly comparing two groups of news, we initially compare the metadata between them. This process helps us identify the top k news groups in the historical record that share similar metadata attributes. Subsequently, we assess the similarity between these top k groups and the news groups from the target trading date.
Figure 4. This diagram illustrates the process by which the News Analyzer assesses similarities across various news collections. As news items pass through the enrichment pipeline, they are tagged with multiple attributes. Each circle in the figure represents one of these attributes. To identify similar news collections from historical data, the analyzer starts by comparing these attributes and filtering out certain ones. Subsequently, it evaluates the similarity among the remaining attributes to determine the connections between different news collections.
@@
In this way, it is efficient to identify the two groups that are similar, then we can make an inference for the target trading date about how these results could impact the following stock market based on the historically similar news and its corresponding market reactions.
3.4.3. Time Decay Hyper-parameter and Dynamic Moving Time Window Training
@@
Figure 5. This figure illustrates the methodology for modeling the impact of earnings conference calls on the stock market when earnings conference calls are not available. It uses a curve to represent the time decay effect of an earnings conference call’s influence over time. When an earnings conference call is initially issued, its impact on the market is at its peak. Over time, this influence gradually diminishes.
@@
Figure 6. This figure provides a visual representation of the Rolling Window methodology in action. By establishing a fixed window length, the approach systematically progresses the training set forward, day by day, in alignment with the passage of time.
4. Conclusion
In this study, we explored the utilization of LLMs in predicting financial risks and introduced the RiskLabs framework. This innovative framework employs LLMs to systematically organize and analyze diverse financial data types and sources, augmenting deep learning models in financial risk prediction. Central to RiskLabs are specialized modules: the Earnings Conference Call Encoder, the Time-Series Encoder, and the News-Market Reactions Encoder, each designed to process specific financial data. These encoders collectively facilitate the merging of various data features for robust multi-task financial risk forecasting. The empirical findings from our study led to several key insights: 1) The RiskLabs framework demonstrates a high efficacy in predicting financial risks, confirming its potential as a valuable tool in this domain. 2) While LLMs in isolation may not yield effective financial risk predictions, their strategic application in processing relevant financial data significantly enhances the predictive power of deep learning models. 3) Ablation studies further affirm that each individual module of RiskLabs meaningfully contributes to the accuracy of the final risk predictions. Overall, our research underscores the transformative potential of LLMs in financial risk assessment, marking a significant step forward in the application of AI in finance.
Appendix A Experiment Details
A.1. Dataset
A.2. Baseline Setup
@@
  • Classical Methods: The analysis incorporates the GARCH model, a classical auto-regressive model for predicting volatility, as well as its various derivatives (Franses and Van Dijk, 1996; Kim and Won, 2018). These models are widely recognized and commonly employed in the realm of volatility prediction. Primarily, they are tailored for short-term volatility forecasting and may not perform as efficiently in predicting average volatility over an extended period (e.g., n-day volatility).
  • LSTM (Gers et al., 2000): Long Short-Term Memory Networks (LSTMs) are a popular choice for financial time series prediction due to their efficacy in handling sequential data. In the context of volatility prediction, we select a straightforward LSTM model as a benchmark.
  • MT-LSTM+ATT (Luong et al., 2015): merges the prediction of average n-day volatility with the forecasting of single-day volatility, using attention-enhanced LSTM as the foundational learning models.
  • HAN (Glove): This baseline implements a Hierarchical Attention Network that incorporates dual-layered attention mechanisms at both the word and sentence levels. Initially, every word within a sentence is transformed into a word embedding through the pre-trained Glove 300-dimensional embeddings. Subsequently, these embedded sentences are processed by a Bi-GRU encoder (Chung et al., 2014), and in parallel, another Bi-GRU encoder is utilized to formulate a representation of each document as a series of sentences. This representation of the document is then fed into the concluding regression layer to generate predictions.
  • MRDM (Qin and Yang, 2019): The MRDM model first introduced a multi-modal deep regression approach for volatility prediction tasks. It utilizes pre-trained GloVe embeddings and bespoke acoustic features, which are processed through individual BiLSTMs to generate uni-modal contextual embeddings. These embeddings are subsequently merged and input into a two-layer dense network for further processing.
  • HTML (Yang et al., 2020): This work presented a state-of-the-art model that employs WWM-BERT for text token encoding. Similar to MDRM, HTML also leverages the same audio features. These unimodal features are then combined and processed through a sentence-level transformer, resulting in multimodal representations for each call.
@@
Appendix B Detailed Experiment Results
B.1. Detailed analysis of VaR values